OpenStudy (anonymous):

Will someone show their work on this problem. 2x/x^2-4 = 4/x^2-4 - 3/x+2 I got No Solution which is the correct answer but I want to make sure I did the problem correctly.

OpenStudy (anonymous):

Is it \[ \frac{2 x}{x^2-4}=\frac{4}{x^2-4}-\frac{3}{x+2 } \]

OpenStudy (anonymous):

Yes and how are you doing that so neatly?

OpenStudy (anonymous):

Uusing LaTeX

OpenStudy (anonymous):

Your equation does not have a solution.

OpenStudy (anonymous):

did you get 15x=0? which is no solution

OpenStudy (precal):